# Singular meanders

### v1

@@ Introduction
A meander is a configuration of a pair of simple curves in a disk. Examples of meanders are shown in Fig. 1. The problem of counting meanders was initially formulated by V. Arnol’d in [A88]; a similar problem of counting closed meanders (under the name of planar permutation) was formulated by P. Rosenstiehl in [R84]. Meanders appear in different areas of mathematics, and we refer readers to the survey [Z23] for both connections and historical context. At present, the meander problem is far from being solved: counts are known only for small numbers of intersections; no subexponential-time algorithms for computing these numbers are known; and the asymptotic growth rate remains conjectural.
@@
In a recent paper [B24], we introduced the factorization of meanders into prime components, and in this context more general objects naturally arise — singular meanders. These are meanders in which not only transverse intersections but also tangency points are allowed (see the example in Fig. 2). Although singular meanders have not been studied before, preliminary enumerative data suggest links with other combinatorial families (some of which were already mentioned in [B24]). In the present note, we provide a preliminary analysis of singular meanders: we discuss some of their properties, and provide some numerical observations.
@@ 1. Definitions
In this section, we present all notation related to singular meanders that we will need later (we refer the reader to [B24, Section 2] for a more detailed discussion). In Subsection 1.1, we define closed singular meanders.
@@
The numbers of singular meanders, iterated snakes and irreducible singular meanders of small total order can be found in [Bcode].
@@ 1.1. Closed singular meanders
As in the case of classical meanders, one can introduce singular versions of other related objects: semimeanders (configurations of a ray and a curve, see [DFGG95]) and stamp foldings (configurations of a segment and a curve, see [koehler1968folding]). One could also consider more general objects allowing both singular intersection points and multiple connected components.
@@ 3.1. Singular meanders with a small number of transverse intersections
The sequence $\{\mathcal{M}_{1,k}\}_{k\geqslant 0}$ coincides with the OEIS sequence A082590 [oeis], which counts the number of $k$-letter words over the alphabet $\{1,2,3,4\}$ having as many occurrences of the substring (consecutive subword) $[1,2]$ as of $[1,3]$. An explicit bijection between equivalence classes of singular meanders with a single transverse intersection and such words will be given in a subsequent paper (together with several other interpretations of singular meanders with a single transverse intersection).
@@ 3.2. Iterated snakes
be the generating function for iterated snakes. Then $F(x,t)$ can be found from the following equation (see [B24]):
(1) $v^{3}+xv^{2}+2v-4(t-1)^{2}(v+x)=0\quad\text{where}\quad v=2F(x,t)+x+2.$
Some subsequences of the array $\left\{\mathcal{M}^{(IS)}_{n,k}\right\}_{n,k\geq 0}$ coincide with other combinatorial sequences that can be found in the OEIS [oeis] [^1] For an exact match, singular meanders of total order 1 should also be considered as snakes.:
@@
The first three sequences coincide since their first terms agree and they satisfy the same recurrence relation (which follows from the Eq. (1)). The sequence A007165 counts the number of $P$-graphs, which are just graph-theoretic reformulations of iterated snakes; see [R86].
@@ 3.3. Irreducible meanders
    $\mathcal{M}^{(Ir)}_{2n-1,4}=n\bigl(\varphi(n+4)-2\bigr)$;
@@
    $\sum_{k\geqslant 0}\mathcal{M}^{(Ir)}_{1,k}t^{k}=\frac{1}{8}\sqrt{\left(\frac{1-3t}{1+t}\right)^{4}-1}$;
@@
    $\sum_{k\geqslant 0}\mathcal{M}^{(Ir)}_{2,k}t^{k}=\frac{2t-1+\sqrt{\frac{1-3t}{1+t}}}{\sqrt{(1-3t)(1+t)^{5}}}$.
@@
Cases (1)–(3) were proved in [B24]. Case (4) follows from case (3) and Theorem 3. Analogously, case (6) follows from case (5) and Theorem 3. The only thing that remains to prove is case (5).
To prove case (5), we introduce two additional classes of singular meanders. A singular meander is said to be arborescent [^2] We use this term because the poset of its submeanders (see [B24]) is a tree. if it can be obtained from an irreducible singular meander by finitely many insertions of irreducible singular meanders. The generating function for arborescent singular meanders of order $(1,k)$ can be easily expressed in terms of the generating function for irreducible singular meanders of order $(1,k)$. Let $\mathcal{M}^{(IIr)}_{n,k}$ be the number of non-equivalent arborescent singular meanders of order $(n,k)$, and set
@@
$G(t)$ $:=\sum_{k\geqslant 0}\mathcal{M}^{(IIr)}_{1,k}t^{k}=\sum_{r\geqslant 0}F(t)^{r}=\frac{1}{1-F(t)}.$
@@
From Theorem 1 it follows that each singular meander of order $(1,k)$ is uniquely obtained via a sequential insertion of B-meanders and iterated snakes, where iterated snakes are not inserted into iterated snakes and B-meanders are not inserted into B-meanders. At the level of generating functions this leads to the following equation:
(2) $C(t)=\frac{A(t)+B(t)+2A(t)B(t)}{1-A(t)B(t)},$

### v2

@@ Introduction
A meander is a configuration of a pair of simple curves in a disk. Examples of meanders are shown in Fig. 1. The problem of counting meanders was initially formulated by V. Arnol’d in [ARN88]; a similar problem of counting closed meanders (under the name of planar permutation) was formulated by P. Rosenstiehl in [ROS84]. Meanders appear in different areas of mathematics, and we refer readers to the survey [ZVO23] for both connections and historical context. At present, the meander problem is far from being solved: counts are known only for small numbers of intersections; no subexponential-time algorithms for computing these numbers are known; and the asymptotic growth rate remains conjectural.
@@
In a recent paper [BEL24], we introduced the factorization of meanders into prime components, and in this context more general objects naturally arise — singular meanders. These are meanders in which not only transverse intersections but also tangency points are allowed (see the example in Fig. 2). Although singular meanders have not been studied before, preliminary enumerative data suggest links with other combinatorial families (some of which were already mentioned in [BEL24]). In the present note, we provide a preliminary analysis of singular meanders: we discuss some of their properties, and provide some numerical observations.
@@ 1. Definitions
In this section, we present all notation related to singular meanders that we will need later (we refer the reader to [BEL24, Section 2] for a more detailed discussion). In Subsection 1.1, we define closed singular meanders.
@@
The numbers of singular meanders, iterated snakes and irreducible singular meanders of small total order can be found in [3].
@@ 1.1. Closed singular meanders
As in the case of classical meanders, one can introduce singular versions of other related objects: semimeanders (configurations of a ray and a curve, see [DGG97]) and stamp foldings (configurations of a segment and a curve, see [KOE68]). One could also consider more general objects allowing both singular intersection points and multiple connected components.
@@ 3.1. Singular meanders with a small number of transverse intersections
The sequence $\{\mathcal{M}_{1,k}\}_{k\geqslant 0}$ coincides with the OEIS sequence A082590 [6], which counts the number of $k$-letter words over the alphabet $\{1,2,3,4\}$ having as many occurrences of the substring (consecutive subword) $[1,2]$ as of $[1,3]$. An explicit bijection between equivalence classes of singular meanders with a single transverse intersection and such words will be given in a subsequent paper (together with several other interpretations of singular meanders with a single transverse intersection).
@@ 3.2. Iterated snakes
be the generating function for iterated snakes. In this subsection singular meanders of total order $1$ are regarded as snakes, in accordance with [BEL24]. Set
$U(x,t):=1+F(x,t).$
Then $U(0,0)=1$, and $U(x,t)$ is the unique formal power series satisfying the equation
(1) $(1-t)U^{3}-\bigl(4(1-t)^{2}-2(1-t)x+1\bigr)U^{2}+(6(1-t)-x)U-2=0.$
Some subsequences of the array $\left\{\mathcal{M}^{(IS)}_{n,k}\right\}_{n,k\geq 0}$ coincide with other combinatorial sequences that can be found in the OEIS [6]:
@@
The first three identifications follow since the corresponding subsequences have the same initial terms and satisfy the same recurrence relation (which follows from Eq. (1)). The sequence A007165 counts the number of $P$-graphs, which are just graph-theoretic reformulations of iterated snakes; see [REA86].
@@ 3.3. Irreducible meanders
    $\mathcal{M}^{(Ir)}_{2n+1,4}=(n+1)\bigl(\varphi(n+5)-2\bigr)$;
@@
    $\sum_{k\geqslant 0}\mathcal{M}^{(Ir)}_{1,k}t^{k}=\frac{1}{8}\left(\sqrt{\frac{1-3t}{1+t}}-1\right)^{4}$;
@@
    $\sum_{k\geqslant 0}\mathcal{M}^{(Ir)}_{2,k}t^{k}=\frac{1-2t-\sqrt{\frac{1-3t}{1+t}}}{\sqrt{(1-3t)(1+t)^{5}}}$.
@@
Cases (1)–(3) were proved in [BEL24]. Case (4) follows from case (3) and Theorem 3. Analogously, case (6) follows from case (5) and Theorem 3. The only thing that remains to prove is case (5).
To prove case (5), we introduce two additional classes of singular meanders. A singular meander is said to be arborescent [^1] We use this term because the poset of its submeanders (see [BEL24]) is a tree. if it can be obtained from an irreducible singular meander by finitely many insertions of irreducible singular meanders. The generating function for arborescent singular meanders of order $(1,k)$ can be easily expressed in terms of the generating function for irreducible singular meanders of order $(1,k)$. Let $\mathcal{M}^{(IIr)}_{n,k}$ be the number of non-equivalent arborescent singular meanders of order $(n,k)$, and set
@@
$G(t)$ $:=\sum_{k\geqslant 0}\mathcal{M}^{(IIr)}_{1,k}t^{k}=\sum_{r\geqslant 1}F(t)^{r}=\frac{F(t)}{1-F(t)}.$
@@
From Theorem 1 it follows that each singular meander of order $(1,k)$ is uniquely obtained via a sequential insertion of B-meanders and iterated snakes(excluding snake of order $(1,0)$), where iterated snakes are not inserted into iterated snakes and B-meanders are not inserted into B-meanders. Since $A(t)$ includes the trivial iterated snake of order $(1,0)$, the generating function for inserted iterated snakes is A(t)1A(t)-1. Therefore
(2) $C(t)=A(t)+\frac{A(t)^{2}B(t)}{1-\left(A(t)-1\right)B(t)},$